# New and updated timing models for seven young energetic X-ray pulsars, including the Big Glitcher PSR J05376910

### v1

@@ Abstract
pulsars: general – pulsars: individual: PSR J00587218, PSR J05376910, PSR B054069, PSR J1412+7922, PSR J18111925, PSR J18131749, PSR J18490001– stars: neutron – X-rays: stars1 Introduction
@@
An outline of the paper is as follows. A brief summary of each pulsar is given next in Section 1.1. Section 2 describes the Chandra, NICER, and NuSTAR data analyzed in this work and their processing. Section 3 presents our results for each of the seven pulsars, including their timing models. Section 4 summarizes the work presented here.
@@ 1.1 Summary of pulsars
PSR J1412+7922 (also known as Calvera) is a high Galactic latitude pulsar (Zane et al., 2011; Halpern et al., 2013) residing within a diffuse radio ring that seems to be a supernova remnant (Arias et al., 2022; Rigoselli et al., 2024; Greco et al., 2025). As a result, PSR J1412+7922 appears to be a member of the class of neutron stars known as central compact objects (CCOs; see De Luca 2017, for review). If the distance of determined from modeling the pulsar’s X-ray spectrum (Mereghetti et al., 2021) and the association between the pulsar and supernova remnant candidate are correct, then the age of the system is (Greco et al., 2025). Monitoring observations of PSR J1412+7922 during the first year of NICER in 2017–2018 yielded its phase-connected timing model (Bogdanov et al., 2019), and subsequent works gradually extended the timing model to over six years using NICER data through 2023 November (Mereghetti et al., 2021; Ho et al., 2022; Rigoselli et al., 2024). Timing analysis by Ho et al. (2022) also yielded a pulsar position that is inconsistent with the position and proper motion measurements made by Halpern & Gotthelf (2015) using Chandra imaging data from 2007 and 2014. Instead, the timing position and 2014 imaging position imply a proper motion of and . A revised measurement of the position and proper motion by Rigoselli et al. (2024) using newer Chandra imaging data from 2024 found values that are much more consistent with the timing results of Ho et al. (2022) (see also Section 3.4), including a much more significant measurement of and .
@@
PSR J18131749 is a young highly energetic pulsar that produces a pulsar wind nebula and is associated with the gamma-ray/TeV source IGR J181351751/HESS J1813178. The pulsar’s proper motion from the center of the young supernova remnant G12.820.02 implies an age of 1000–2200 yr (Dzib & Rodríguez, 2021). PSR J18131749 has a broad single peak X-ray pulse profile and high X-ray pulsed fraction (Gotthelf & Halpern, 2009; Halpern et al., 2012; Kuiper & Hermsen, 2015; Ho et al., 2020a; Takata et al., 2024). Even though the pulsar has also been detected in radio, its radio pulses suffer from extremely high scattering (Camilo et al., 2021).
@@ 2.1 Chandra data
Chandra observed PSR J05376910 twice in 2002 using the ACIS-S detector on August 23 and October 29 (ObsID 2783) for a total of about 48 ks. Chandra observed PSR J18131749 six times using the ACIS-S detector in continuous clocking (CC) mode on 2012 February 12 (ObsID 12549), 2021 February 10 (ObsID 23545), 2021 June 23 (ObsID 23546), 2024 April 3 (ObsID 28355), 2024 October 7 (ObsID 28356), and 2025 March 23 (ObsID 28357) for about 20 ks on each date (see Table 2). We reprocess these data following the standard procedure using chandra_repro of the Chandra Interactive Analysis of Observations (CIAO) package version 4.17 and Calibration Database (CALDB) 4.12.0 (Fruscione et al., 2006). For data on PSR J05376910, we use specextract to extract source events from a 1″ radius circle centered on the pulsar position and with a background region consisting of a 3″10″ rectangular box centered on the pulsar but excluding the source extraction circle (see Figure 4 of Chen et al. 2006). For data on PSR J18131749, we extract source events from the one-dimensional CC image along a 15 length centered on the pulsar position and with a background region 11″ in length and about 10″ from the source region; we also consider a larger 3″ extraction length to collect more photons for spectral analysis (see Section 3.6), although this gives somewhat weaker pulse detections. Spectra are combined using combine_spectra and binned using dmgroup with a minimum of 25 counts per bin.
@@ 2.2 NICER data
We process and filter NICER data using nicerl2 of HEASoft 6.22–6.35.2 (HEASARC, 2014) and NICERDAS 2018-03-01_V003–2025-06-11_V014. While NICER is sensitive to 0.25–12 keV photons, we extract only events within a particular energy range for each pulsar to optimize pulsation searches (1–7 keV for PSR J05376910, 0.37–1.97 keV for PSR J1412+7922, and 1.89–6 keV for PSR J18490001). We ignore time intervals of enhanced background affecting all detectors by constructing a light curve binned at 16 s and removing intervals strongly contaminated by background flaring when the count rate exceeds a threshold value that is different for each pulsar (10 c s-1 for PSR J05376910, 4.5 c s-1 for PSR J1412+7922, and 5 c s-1 for PSR J18490001). Using these filtering criteria, we obtain clean data for pulse timing analysis. We use barycorr to transform between TT, used for event time stamps, and TDB and to account for effects of satellite motion with respect to the barycenter. In all timing analyses performed here unless otherwise noted (in particular, Sections 3.4 and 3.7), source positions are held fixed at the values given in the corresponding tables below, along with our adopted solar system ephemeris. As with analysis of Chandra data, we use PRESTO and prepfold to perform a pulsation search and to determine the spin frequency of each pulsar. A template pulse profile for each pulsar is produced by fitting a set of NICER pulse profiles with a Gaussian shape; this template is then used to determine the TOA of each observation following the unbinned maximum likelihood technique described in Ray et al. (2011). Unless otherwise noted below, we use TEMPO2 (Hobbs et al., 2006) to fit TOAs with a timing model and to measure glitch parameters. In particular, a pulsar’s rotation phase is fit to a reference time based on a truncated Taylor series of the spin frequency and its time derivatives,
@@ 2.3 NuSTAR data
NuSTAR observed PSR J05376910 on 2016 October 17 (ObsID 40201014002) for 105 ks. We reprocess this data using nupipeline and nuproducts of HEASoft 6.35.2 and NUSTARDAS 2025-03-11_v2.1.5 and barycenter events with clock correction 20100101v211 and the pulsar’s position. The source extraction region is a circle of 60″ radius, and the background region is a circle of the same size and placed in a field with low counts. For timing analysis, we use xselect to choose only events at 3–20 keV. We use PRESTO and prepfold to perform pulsation searches on the merged FPMA and FPMB event list and to determine the spin frequency. To extract pulsed spectra, we first use photonphase in PINT (Luo et al., 2021) and the spin frequency to determine the rotation phase of each photon in the data. We then use maketime to determine GTIs for on-pulse emission and off-pulse emission. Finally, we run nuproducts with these GTIs to produce on and off-pulse spectra.
@@ 3.2 PSR J05376910
Because of the fast spin of PSR J05376910 and that its spin pulsations are only detectable in X-ray, few instruments have been able to track its timing properties over long periods of time. RXTE performed high-cadence observations of PSR J05376910 from 1999 to 2011, and these data yielded phase-connected timing models during this 13-year period (Antonopoulou et al., 2018; Ferdman et al., 2018). NICER started regular monitoring observations of PSR J05376910 soon after its launch in mid-2017. Phase-connected timing models and 8 glitches measured with NICER data through 2020 April were presented in Ho et al. (2020b). Subsequent timing models through 2020 October, including 3 further glitches, and then through 2022 February, including 4 more glitches, were reported in Abbott et al. (2021a) and Ho et al. (2022), respectively. Here we report timing models using the remaining NICER data from 2022 March through 2025 June, with 8 new glitches detected during this period (first 6 of which are also reported in Zubieta et al. 2026). Tables 4 and 5 show timing models and glitch parameters derived from the complete NICER dataset from 2017 to 2025, including those given in our previous works above. We use our previous naming convention, where each segment is separated by a glitch and is labeled by glitch number, with segment 1 occurring after glitch 1, which is the first NICER-detected glitch. Figure 3 shows residuals of our best-fit timing models to all the NICER data, Figure 4 plots the spin-down rate measured by RXTE and NICER since 1999, and Figure 5 displays glitch parameters and for the 23 glitches measured using NICER. A best-fit linear decline for the long-term spin-down rate yields and a braking index calculated using and from around the middle of the 25 years of RXTE and NICER observations. For a few glitches, an exponential recovery term is included in the glitch model [see equation (2)], but the presence of the recovery is not considered when characterising the intervals between glitches. A detailed study of exponential recoveries of glitches of PSR J05376910 is conducted in Zubieta et al. (2026).
@@
Here we provide a more straightforward measurement of the pulsed spectrum at X-ray energies 3–79 keV using a 105 ks NuSTAR observation from 2016 (see Table 2). Note that Bamba et al. (2022) analyzed this same NuSTAR data, alongside Suzaku data, to determine the total (pulsar+pulsar wind nebula) spectrum at 1.5–79 keV but did not perform timing analysis to extract the pulsed spectrum of PSR J05376910. First, we follow the procedure described in Section 2.3 and measure a spin frequency at MJD 57678.64, where the uncertainty in the last digit is given in parentheses. Figure 8 shows the resulting 3–20 and 20 keV pulse profiles when the data are folded at this frequency, alongside the 1–7 keV pulse profile measured using NICER data. This spin frequency is higher than the frequency predicted assuming the approximate pulsar spin-down rate during the time since the last RXTE data five years prior to the NuSTAR observation (see Figure 4). Such a deviation can be expected, given that PSR J05376910 likely underwent spin-up glitches during this time. Analogously, the NuSTAR frequency is lower than the frequency extrapolated from the first NICER data. This is consistent with a few spin-up glitches during the intervening year between NuSTAR and NICER observations.
@@ 3.3 PSR B054069
NICER observations of PSR B054069 began on 2017 July 25. Previous works reported timing models using NICER data from a relatively short time range of 2023 January 19 to May 3, supplemented by IXPE data (Xie et al., 2024), and from 2022 November 18 to 2023 December 25 (Tuo et al., 2024). In Espinoza et al. (2024), we presented results from our timing models using NICER data from the start of its observations on 2017 July 25 through 2024 July 29, supplemented by IXPE and Swift data. Here we provide the timing models and update these models to include all the effective NICER data through 2025 May 20. Figure 10 shows timing residuals of the TOAs used to obtain our best-fit timing models, which are given in Table 6. Here the timing model segments for PSR B054069 are not separated by glitches, in contrast to the glitch-separated segments in the timing model for, e.g., PSR J05376910 given in Section 3.2. The segments are determined by periods over which a model that includes provides a resonable fit to the data.
@@ 3.5 PSR J18111925
NICER monitoring of PSR J18111925 began with a few observations in 2021 September and then commenced in earnest in 2022 March. Here we report the timing model derived from all NICER data through 2025 May. Analysis of these data suggest PSR J18111925 suffers from timing noise. As a result, we find the TOAs are better modeled by dividing them into two epochs: TOAs obtained before the period in late 2023 when PSR J18111925 cannot be observed due to its proximity to the Sun in the sky and TOAs obtained after this period in early 2024, i.e., before about MJD 60227 and after about MJD 60390. Figure 13 shows residuals of our best-fit timing models for these two epochs, which are given in Table 8, and Figure 14 shows the pulse profile at 2.5–8 keV from the combined NICER observations. We note that we are able to construct a single timing model for the entire period of NICER observations, but such a model requires two additional spin terms (fourth and fifth time derivatives of ) and yields a best-fit with RMS of 2.15 ms and .
@@ 3.7 PSR J18490001
Figure 16 shows timing residuals of the TOAs used to obtain our best-fit timing model, which is given in middle column of Table 10. We find that our position from timing is consistent with the Chandra HRC-S imaging position from Kuiper & Hermsen (2015) of R.A. and decl. at MJD 55885, which has a 90 percent confidence level uncertainty of 06. It is noteworthy that our timing position, with an uncertainty of 011, is now more accurate than the one from imaging.

### v2

@@ Abstract
stars: neutron – pulsars: general – pulsars: individual: PSR J00587218, PSR J05376910, PSR B054069, PSR J1412+7922, PSR J18111925, PSR J18131749, PSR J18490001– X-rays: stars1 Introduction
@@
An outline of the paper is as follows. A brief summary of each pulsar is given next in Section 1.1. Section 2 describes the Chandra, NICER, and NuSTAR data analysed in this work and their processing. Section 3 presents our results for each of the seven pulsars, including their timing models. Section 4 summarizes the work presented here.
@@ 1.1 Summary of pulsars
PSR J1412+7922 (also known as Calvera) is a high Galactic latitude pulsar (Zane et al., 2011; Halpern et al., 2013) residing within a diffuse radio ring that seems to be a supernova remnant (Arias et al., 2022; Rigoselli et al., 2024; Greco et al., 2025). As a result, PSR J1412+7922 appears to be a member of the class of neutron stars known as central compact objects (see De Luca 2017, for review). If the distance of determined from modeling the pulsar’s X-ray spectrum (Mereghetti et al., 2021) and the association between the pulsar and supernova remnant candidate are correct, then the age of the system is (Greco et al., 2025). Monitoring observations of PSR J1412+7922 during the first year of NICER in 2017–2018 yielded its phase-connected timing model (Bogdanov et al., 2019), and subsequent works gradually extended the timing model to over six years using NICER data through 2023 November (Mereghetti et al., 2021; Ho et al., 2022; Rigoselli et al., 2024). Timing analysis by Ho et al. (2022) also yielded a pulsar position that is inconsistent with the position and proper motion measurements made by Halpern & Gotthelf (2015) using Chandra imaging data from 2007 and 2014. Instead, the timing position and 2014 imaging position imply a proper motion of and . A revised measurement of the position and proper motion by Rigoselli et al. (2024) using newer Chandra imaging data from 2024 found values that are much more consistent with the timing results of Ho et al. (2022) (see also Section 3.4), including a much more significant measurement of and .
@@
PSR J18131749 is a young highly energetic pulsar that produces a pulsar wind nebula and is associated with the gamma-ray/TeV source IGR J181351751/HESS J1813178. The pulsar’s proper motion from the centre of the young supernova remnant G12.820.02 implies an age of 1000–2200 yr (Dzib & Rodríguez, 2021). PSR J18131749 has a broad single peak X-ray pulse profile and high X-ray pulsed fraction (Gotthelf & Halpern, 2009; Halpern et al., 2012; Kuiper & Hermsen, 2015; Ho et al., 2020a; Takata et al., 2024). Even though the pulsar has also been detected in radio, its radio pulses suffer from extremely high scattering (Camilo et al., 2021).
@@ 2.1 Chandra data
Chandra observed PSR J05376910 twice in 2002 using the ACIS-S detector on August 23 and October 29 (ObsID 2783) for a total of about 48 ks. Chandra observed PSR J18131749 six times using the ACIS-S detector in continuous clocking (CC) mode on 2012 February 12 (ObsID 12549), 2021 February 10 (ObsID 23545), 2021 June 23 (ObsID 23546), 2024 April 3 (ObsID 28355), 2024 October 7 (ObsID 28356), and 2025 March 23 (ObsID 28357) for about 20 ks on each date (see Table 2). We reprocess these data following the standard procedure using chandra_repro of the Chandra Interactive Analysis of Observations (CIAO) package version 4.17 and Calibration Database (CALDB) 4.12.0 (Fruscione et al., 2006). For data on PSR J05376910, we use specextract to extract source events from a 1 arcsec radius circle centred on the pulsar position and with a background region consisting of a 3 arcsec10 arcsec rectangular box centred on the pulsar but excluding the source extraction circle (see Figure 4 of Chen et al. 2006). For data on PSR J18131749, we extract source events from the one-dimensional CC image along a 1.5 arcsec length centred on the pulsar position and with a background region 11 arcsec in length and about 10 arcsec from the source region; we also consider a larger 3 arcsec extraction length to collect more photons for spectral analysis (see Section 3.6), although this gives somewhat weaker pulse detections. Spectra are combined using combine_spectra and binned using dmgroup with a minimum of 25 counts per bin.
@@ 2.2 NICER data
We process and filter NICER data using nicerl2 of HEASoft 6.22–6.35.2 (HEASARC, 2014) and NICERDAS 2018-03-01_V003–2025-06-11_V014. While NICER is sensitive to 0.25–12 keV photons, we extract only events within a particular energy range for each pulsar to optimize pulsation searches (1–7 keV for PSR J05376910, 0.37–1.97 keV for PSR J1412+7922, and 1.89–6 keV for PSR J18490001). We ignore time intervals of enhanced background affecting all detectors by constructing a light curve binned at 16 s and removing intervals strongly contaminated by background flaring when the count rate exceeds a threshold value that is different for each pulsar (10 c s-1 for PSR J05376910, 4.5 c s-1 for PSR J1412+7922, and 5 c s-1 for PSR J18490001). Using these filtering criteria, we obtain clean data for pulse timing analysis. We use barycorr to transform between TT, used for event time stamps, and TDB and to account for effects of satellite motion with respect to the barycentre. In all timing analyses performed here unless otherwise noted (in particular, Sections 3.4 and 3.7), source positions are held fixed at the values given in the corresponding tables below, along with our adopted Solar system ephemeris. As with analysis of Chandra data, we use PRESTO and prepfold to perform a pulsation search and to determine the spin frequency of each pulsar. A template pulse profile for each pulsar is produced by fitting a set of NICER pulse profiles with a Gaussian shape; this template is then used to determine the TOA of each observation following the unbinned maximum likelihood technique described in Ray et al. (2011). Unless otherwise noted below, we use TEMPO2 (Hobbs et al., 2006) to fit TOAs with a timing model and to measure glitch parameters. In particular, a pulsar’s rotation phase is fit to a reference time based on a truncated Taylor series of the spin frequency and its time derivatives,
@@ 2.3 NuSTAR data
NuSTAR observed PSR J05376910 on 2016 October 17 (ObsID 40201014002) for 105 ks. We reprocess this data using nupipeline and nuproducts of HEASoft 6.35.2 and NUSTARDAS 2025-03-11_v2.1.5 and barycentre events with clock correction 20100101v211 and the pulsar’s position. The source extraction region is a circle of 60 arcsec radius, and the background region is a circle of the same size and placed in a field with low counts. For timing analysis, we use xselect to choose only events at 3–20 keV. We use PRESTO and prepfold to perform pulsation searches on the merged FPMA and FPMB event list and to determine the spin frequency. To extract pulsed spectra, we first use photonphase in PINT (Luo et al., 2021) and the spin frequency to determine the rotation phase of each photon in the data. We then use maketime to determine GTIs for on-pulse emission and off-pulse emission. Finally, we run nuproducts with these GTIs to produce on and off-pulse spectra.
@@ 3.2 PSR J05376910
Because of the fast spin of PSR J05376910 and that its spin pulsations are only detectable in X-ray, few instruments have been able to track its timing properties over long periods of time. RXTE performed high-cadence observations of PSR J05376910 from 1999 to 2011, and these data yielded phase-connected timing models during this 13-year period (Antonopoulou et al., 2018; Ferdman et al., 2018). NICER started regular monitoring observations of PSR J05376910 soon after its launch in mid-2017. Phase-connected timing models and 8 glitches measured with NICER data through 2020 April were presented in Ho et al. (2020b). Subsequent timing models through 2020 October, including 3 further glitches, and then through 2022 February, including 4 more glitches, were reported in Abbott et al. (2021a) and Ho et al. (2022), respectively. Here we report timing models using the remaining NICER data from 2022 March through 2025 June, with 8 new glitches detected during this period (first 6 of which are also reported in Zubieta et al. 2026). Tables 4 and 5 show timing models and glitch parameters derived from the complete NICER dataset from 2017 to 2025, including those given in our previous works above. We use our previous naming convention, where each segment is separated by a glitch and is labelled by glitch number, with segment 1 occurring after glitch 1, which is the first NICER-detected glitch. Figure 3 shows residuals of our best-fit timing models to all the NICER data, Figure 4 plots the spin-down rate measured by RXTE and NICER since 1999, and Figure 5 displays glitch parameters and for the 23 glitches measured using NICER. A best-fit linear decline for the long-term spin-down rate yields and a braking index calculated using and from around the middle of the 25 years of RXTE and NICER observations. For a few glitches, an exponential recovery term is included in the glitch model [see equation (2)], but the presence of the recovery is not considered when characterizing the intervals between glitches. A detailed study of exponential recoveries of glitches of PSR J05376910 is conducted in Zubieta et al. (2026).
@@
Here we provide a more straightforward measurement of the pulsed spectrum at X-ray energies 3–79 keV using a 105 ks NuSTAR observation from 2016 (see Table 2). Note that Bamba et al. (2022) analysed this same NuSTAR data, alongside Suzaku data, to determine the total (pulsar+pulsar wind nebula) spectrum at 1.5–79 keV but did not perform timing analysis to extract the pulsed spectrum of PSR J05376910. First, we follow the procedure described in Section 2.3 and measure a spin frequency at MJD 57678.64, where the uncertainty in the last digit is given in parentheses. Figure 8 shows the resulting 3–20 and 20 keV pulse profiles when the data are folded at this frequency, alongside the 1–7 keV pulse profile measured using NICER data. This spin frequency is higher than the frequency predicted assuming the approximate pulsar spin-down rate during the time since the last RXTE data five years prior to the NuSTAR observation (see Figure 4). Such a deviation can be expected, given that PSR J05376910 likely underwent spin-up glitches during this time. Analogously, the NuSTAR frequency is lower than the frequency extrapolated from the first NICER data. This is consistent with a few spin-up glitches during the intervening year between NuSTAR and NICER observations.
@@ 3.3 PSR B054069
NICER observations of PSR B054069 began on 2017 July 25. Previous works reported timing models using NICER data from a relatively short time range of 2023 January 19 to May 3, supplemented by IXPE data (Xie et al., 2024), and from 2022 November 18 to 2023 December 25 (Tuo et al., 2024). In Espinoza et al. (2024), we presented results from our timing models using NICER data from the start of its observations on 2017 July 25 through 2024 July 29, supplemented by IXPE and Swift data. Here we provide the timing models and update these models to include all the effective NICER data through 2025 June 10. Figure 10 shows timing residuals of the TOAs used to obtain our best-fit timing models, which are given in Table 6. Here the timing model segments for PSR B054069 are not separated by glitches, in contrast to the glitch-separated segments in the timing model for, e.g., PSR J05376910 given in Section 3.2. The segments are determined by periods over which a model that includes provides a resonable fit to the data.
@@ 3.5 PSR J18111925
NICER monitoring of PSR J18111925 began with a few observations in 2021 September and then commenced in earnest in 2022 March. Here we report the timing model derived from all NICER data through 2025 May. Analysis of these data suggest PSR J18111925 suffers from timing noise. As a result, we find the TOAs are better modelled by dividing them into two epochs: TOAs obtained before the period in late 2023 when PSR J18111925 cannot be observed due to its proximity to the Sun in the sky and TOAs obtained after this period in early 2024, i.e., before about MJD 60227 and after about MJD 60390. Figure 13 shows residuals of our best-fit timing models for these two epochs, which are given in Table 8, and Figure 14 shows the pulse profile at 2.5–8 keV from the combined NICER observations. We note that we are able to construct a single timing model for the entire period of NICER observations, but such a model requires two additional spin terms (fourth and fifth time derivatives of ) and yields a best-fit with RMS of 2.15 ms and .
@@ 3.7 PSR J18490001
Figure 16 shows timing residuals of the TOAs used to obtain our best-fit timing model, which is given in the middle column of Table 10. We find that our position from timing is consistent with the Chandra HRC-S imaging position from Kuiper & Hermsen (2015) of R.A. and decl. at MJD 55885, which has a 90 percent confidence level uncertainty of 0.6 arcsec. It is noteworthy that our timing position, with an uncertainty of 0.11 arcsec, is now more accurate than the one from imaging.